OpenStudy (anonymous):

Solve the system by the elimination method. Check your work. 3a + 5b - 7 = 0 a - 2b - 4 = 0 {(96/11, -5/11)} {(34/11, -5/11)} {(32/33, 5/11)}

OpenStudy (anonymous):

@kl0723 or & @perl can you help?

OpenStudy (kl0723):

ok, your answer should of the form (x,y) that means one point for x and one for y or in this case (a,b)

OpenStudy (kl0723):

you have two variables and must get rid of one of them... try by multyplying the second equations by -3 to make the a values cancel out :)

OpenStudy (anonymous):

I have no calculator or anything on me. That's why im asking for it :c

OpenStudy (kl0723):

don't really need one :) \[3a + 5b -7 = 0\] \[a - 2b -4 = 0\] whole equation times (3)

OpenStudy (anonymous):

youd have to multiple it by -3 not just 3

OpenStudy (kl0723):

I meant (-3), so you will have now -3a +6b +12 = 0... this is your second equation :)

OpenStudy (anonymous):

multiply*

OpenStudy (perl):

3a + 5b - 7 = 0 a - 2b - 4 = 0 Rewrite in the form ax +by = c 3a + 5b = 7 a - 2b = 4

OpenStudy (perl):

ok so far?

OpenStudy (perl):

i added 7 to both sides of the first equation, and added 4 to both sides of the second equation

OpenStudy (anonymous):

I just guessed on my work its ok.

OpenStudy (kl0723):

don't guess! we're trying to make sure you understand how to get to your answer, if we just give you an answer we're not helping at all

OpenStudy (anonymous):

it was just that question! everything else is graphing.

OpenStudy (perl):

now multiply second equation by -3